# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 3.1 Baseline PG-DPO for the Extended Objective
$\lambda_{k}^{(i)}\approx\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r+\pi_{k}^{(i)}(\mu-r)\bigr{]}\Delta t+\lambda_{k+1}^{(i)}\pi_{k}^{(i)}\sigma\,\Delta W_{k}^{(i)}.$
This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$. To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$). Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of a more complete pathwise recurrence relation (inclusive of all terms affecting $J^{(i)}$) and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
@@
$\lambda_{k}=\lambda_{k+1}+\underbrace{\bigl{(}\lambda_{k+1}[r+\pi_{k}(\mu-r)]+\pi_{k}\,\sigma\,Z_{k}\bigr{)}}_{\approx\partial_{x}\mathcal{H}(t_{k},\dots)\text{ from PMP}}\Delta t-Z_{k}\,\Delta W_{k}.$ (14)
Crucially, the drift term derived through this process corresponds exactly to the derivative of the discrete-time Hamiltonian ($\partial_{x}\mathcal{H}$) from PMP with respect to the state variable $x$. A similar structure holds for the costate vector $\bm{\eta}_{k}$ associated with the state variables $\mathbf{Y}_{k}$.
Therefore, the values $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$ computed by BPTT are indeed pathwise numerical counterparts related to the Pontryagin costates, satisfying the relevant discrete BSDE structure. Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t. states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental. Firstly, they are essential inputs for the Projected PGDPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
@@ 3.2 Projected PG-DPO (P-PGDPO) Algorithm
While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge. The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality. This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly by projecting them onto the manifold of policies satisfying the PMP first-order conditions.
@@
This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment. The procedure consists of two distinct stages, described below using a single level of enumeration:
  1. 1.
    Stage 1: Warm-Up Phase for Costate Estimation
    First, we execute the Baseline PG-DPO process, as detailed in Section 3.1 (which optimizes $\widetilde{J}$), for a predetermined, relatively small number of iterations, $K_{0}$. The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize. We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
  2. 2.
    Stage 2: Analytic Control Deployment
@@
In summary, the Projected PG-DPO (P-PGDPO) approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates. It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP. This strategy can significantly reduce the overall computational burden while yielding high-quality, near-optimal controls. The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
@@ 3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012). This HJB equation exhibits a parabolic form. The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables). While the first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) are crucial for the solution’s properties, they do not affect the parabolicity classification, which is mainly governed by the second-order diffusion terms.
The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022). This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf. Evans, , 2022, Ch. 6-7). Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy. The core idea of Projected PG-DPO (P-PGDPO) involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas. Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the P-PGDPO method.
@@
Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Projected PG-DPO (P-PGDPO) method. The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where P-PGDPO effectively leverages the expected regularity of the Costates. The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
Building upon this expected regularity, a more formal mathematical justification for the P-PGDPO approach can be provided. The objective is to establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with BPTT estimation error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$. This analysis relies on certain standing assumptions concerning parabolic regularity and the FOC gap, which are formally stated as Assumption 6 and 7 in Appendix C. We use $L^{q,p}(Q_{T})$ to denote the standard mixed space-time Lebesgue norm over the domain $Q_{T}=(0,T)\times\Omega\subseteq\mathbb{R}^{d+1}$, as detailed in Appendix C.
The main result of this analysis is encapsulated in the following theorem, the proof of which is detailed in Appendix C along with supporting lemmas and propositions.
Theorem 1 (Policy Gap Bound).
Under Assumptions 6 and 7 (detailed in Appendix C), let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$. Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs. Assume $F$ is locally Lipschitz continuous with constant $C_{F}$. If the BPTT estimation error $\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$ (where $\delta_{\mathrm{BPTT}}$ quantifies the precision of BPTT-derived costate estimates relative to Stage 1 costates), then the policy gap is bounded by:
$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
Here, $C_{\mathrm{tot}}$ is a constant that depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ (derived from Proposition C.2 in Appendix C).
The proof of Theorem 1 follows a logical chain: the FOC-gap $(\leq\varepsilon)$ implies that errors in PDE coefficients and source terms are of order $O(\varepsilon)$. These, in turn, through PDE stability, imply that the error in the value function’s gradient $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$ is also of order $O(\varepsilon)$. Combining this with the BPTT estimation error $\delta_{\mathrm{BPTT}}$ leads to the final policy-gap bound. The detailed arguments, including the necessary lemmas and propositions, are provided as sketches in Appendix C. It is important to note that all constants involved depend quantitatively on model parameters, PDE regularity, and utility specifications. Full rigorous proofs would require meticulous tracking of these dependencies and bounds.
Our formal analysis (detailed in Appendix C) demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin first-order condition (FOC) gap ($\varepsilon$) achieved during the Stage 1 warm-up guarantees that the BPTT-derived costate estimates are $O(\varepsilon)$-close to the true optimal costates (up to the BPTT estimation error $\delta_{\mathrm{BPTT}}$). This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these rapidly stabilizing and sufficiently accurate costate estimates into the analytical PMP formulas yields a near-optimal policy $\widehat{\pi}$ with an error bound of $O(\varepsilon+\delta_{\mathrm{BPTT}})$ (Theorem 1).
@@
Despite these formal hurdles, the fundamental algorithmic rationale underpinning the P-PGDPO approach remains persuasive for broader applicability. Neither Pontryagin’s Maximum Principle (PMP) (Pontryagin, , 2014; Yong and Zhou, , 2012) nor backpropagation-through-time (BPTT) (Goodfellow, , 2016) are inherently limited to uniformly parabolic or affine settings. The P-PGDPO strategy leverages the potentially advantageous decoupling of the problem into: (i) estimating potentially simpler or faster-converging costate functions via BPTT, and (ii) utilizing the robust analytical structure of PMP to construct the control.
This rationale, strongly supported by the numerical success demonstrated even in high dimensions for the affine benchmark case (Section 4), suggests that PG-DPO (especially the P-PGDPO variant) holds promise as a practical tool for a wider class of challenging continuous-time control problems. Nonetheless, confirming its effectiveness and potentially refining the theoretical guarantees for models outside the strictly regular class analyzed here remains an important avenue for future research, likely requiring tailored analysis or different analytical techniques (e.g., focusing on regions predominantly visited by the state process).

### v4

@@ 3.1 Baseline PG-DPO for the Extended Objective
$\mathbb{E}\!\Bigl{[}\lambda_{k+1}^{(i)}(\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)})\Bigm{|}\mathcal{F}_{t_{k}}\Bigr{]}\;=\;\bigl{(}\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}\bigr{)}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})\,\Delta t.$
Proof.
The product inside the expectation is an inner product of vectors. The term involving the adapted part, $\mathbb{E}[\lambda_{k+1}(\dots)|\mathcal{F}_{t_{k}}]$, is zero as it is the expectation of a martingale increment. The non-zero terms arise from the quadratic covariation of the raw martingale components of $\lambda_{k+1}^{(i)}$ with the wealth shock. Using the standard properties $\mathbb{E}[(\Delta\mathbf{W}_{k}^{X})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=I\Delta t$ and $\mathbb{E}[(\Delta\mathbf{W}_{k}^{Y})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=\bm{\rho}_{Y\!X}\Delta t$, where $\bm{\rho}_{Y\!X}$ is the cross-correlation matrix, yields the stated result. ∎
Theorem 3 (BPTT–PMP Correspondence).
Define the effective martingale component as $\mathbf{Z}^{\lambda X}_{k}:=\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}$. The adapted costate process $(\lambda_{k})_{k=0}^{N}$ satisfies the discrete-time backward stochastic difference equation:
@@
$\lambda_{k}$ $=\lambda_{k+1}+\Bigl{[}\lambda_{k+1}\bigl{(}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}+\mathbf{Z}^{\lambda X}_{k}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})+\bigl{(}e^{-\delta t_{k}}U^{\prime}(C_{k})-\lambda_{k+1}\bigr{)}\frac{\partial C_{k}}{\partial X_{k}}\Bigr{]}\Delta t$
$\quad-\widetilde{\mathbf{Z}}^{\lambda X}_{k}\,\Delta\mathbf{W}^{X}_{k}-\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\,\Delta\mathbf{W}^{Y}_{k}.$ (16)
Furthermore, the drift term in the square brackets of Eq. (3) is identical to the partial derivative of the Hamiltonian, $\partial_{x}\mathcal{H}$, evaluated at the corresponding state and control variables.
Proof.
We begin by taking the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}_{t_{k}}]$ of the entire pathwise recurrence in Eq. (1) from Lemma 1. The left-hand side becomes $\lambda_{k}$ by definition. On the right-hand side, the expectations of the drift terms are straightforward. The crucial step is handling the stochastic term $\mathbb{E}[\lambda_{k+1}^{(i)}(\dots)|\mathcal{F}_{t_{k}}]$, for which we substitute the result from Lemma 2. This transforms the stochastic term into the additional drift component involving the effective process $\mathbf{Z}^{\lambda X}_{k}$. After collecting all drift terms and separating the martingale components, we arrive at the discrete-time BSDE presented in Eq. (3). The final step involves a direct algebraic comparison of this drift term with the definition of the Hamiltonian $\mathcal{H}$ from Section 2.2, which confirms their identity. ∎
Remark 1.
The Martingale Representation Theorem, cited from Karatzas and Shreve, (1991), is a continuous-time result. Its application in our discrete-time grid setting is justified as the discrete process can be viewed as a sampling of an underlying continuous process, for which the representation holds on each interval $[t_{k},t_{k+1}]$.
This theorem provides the theoretical guarantee that our “Pontryagin-guided” approach is well-founded. BPTT is not merely a generic optimization heuristic; it is a computational procedure that implicitly solves the necessary optimality conditions of PMP. While the baseline PG-DPO uses this powerful link implicitly to find effective policy gradients, our Projected PG-DPO variant leverages it explicitly by taking the BPTT-computed costates and using them to construct a policy that is structurally consistent with financial theory.
@@ 3.2 Projected PG-DPO (P-PGDPO) Algorithm
While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it may require a large number of iterations for the networks to fully converge. The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative, based on the empirical observation that the costate estimates—$(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$ and their relevant derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$)—often stabilize and become reasonably accurate well before the policy networks reach convergence.
P-PGDPO leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly, by projecting them onto the manifold of policies that satisfy the PMP first-order conditions. In this way, it bypasses the need for fully converged policy networks in order to deploy effective controls.
The procedure unfolds in two distinct stages, as described below using a single-level enumeration:
  1. 1.
    Stage 1: Warm-Up Phase for Costate Estimation
    First, we execute the Baseline PG-DPO procedure (see Section 3.1), which optimizes $\widetilde{J}$, for a predetermined, relatively small number of iterations $K_{0}$. The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize. We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
  2. 2.
    Stage 2: Analytic Control Deployment
@@
In summary, the Projected PG-DPO (P-PGDPO) approach begins with a short warm-up phase (Stage 1), using the baseline PG-DPO procedure to obtain a preliminary policy $(\theta^{*},\phi^{*})$ whose backpropagation-through-time (BPTT) produces reasonably stable costate estimates. In Stage 2, these costate estimates are refined via Monte Carlo averaging at deployment time and then substituted into the analytical expression for the optimal control derived from PMP.
This two-stage strategy can substantially reduce computational costs while delivering high-quality, near-optimal controls. The Monte Carlo averaging step in Stage 2 is essential for obtaining reliable inputs to the PMP formula, mitigating the noise inherent in single-path estimates.
@@ 3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are governed by stochastic differential equations (SDEs), the value function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically satisfies a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012). The HJB equation is parabolic, and the specific nature of its parabolicity is primarily determined by the diffusion coefficients in the underlying SDEs—namely $\bm{\sigma}$ (for asset dynamics) and $\bm{\sigma}_{Y}$ (for state dynamics). While first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) influence the structure of the solution, they do not affect the parabolicity classification, which is governed by the second-order terms.
The behavior of the diffusion matrix $a(t,x,\mathbf{y})$, which multiplies the second-order derivatives in the HJB equation, allows the PDE to be classified into different parabolic types (see, e.g., Evans, , 2022). This classification significantly impacts the regularity (smoothness) of the value function $V$ and its associated costates $\nabla V$ (cf. Evans, , 2022, Chs. 6–7). Costates represent the sensitivity of the value function with respect to states and wealth, and Pontryagin’s Maximum Principle links these costates directly to the optimal policy.
The core idea of the Projected PG-DPO (P-PGDPO) method is to estimate these costates—and their derivatives—via BPTT and substitute them into the analytical expressions for the optimal control provided by PMP. The regularity of costates, governed by the parabolicity class of the HJB equation, directly influences the accuracy of these estimates and thus the effectiveness of P-PGDPO.
@@
The reason this formal justification is crucial lies in a fundamental challenge facing many simulation-based optimization methods. A critical issue in direct policy optimization, including standard reinforcement learning, is that a learned policy can achieve a near-optimal objective value (a small value gap) while being structurally very different from the true optimal policy (a large policy gap). This occurs because the objective function is often locally “flat” near the optimum, meaning first-order policy deviations may only cause second-order utility losses.
Our framework circumvents this issue by shifting focus from the value itself to the satisfaction of Pontryagin’s first-order conditions (FOCs). The success of this approach hinges on a key property of the HJB equations common in finance: their parabolic nature. Parabolic PDEs, akin to the heat equation, enforce a “smoothing” effect, ensuring that a well-behaved value function also has well-behaved derivatives (costates). This is not merely an intuitive notion; it is mathematically formalized by regularity theory in Sobolev spaces, which guarantees that convergence in value implies convergence of its derivatives. This regularity is precisely what allows us to trust the BPTT-derived costates and reconstruct a robust policy via the PMP formula.
Building on this expected regularity, we provide a more formal justification for P-PGDPO. We aim to bound the policy gap, $\|\widehat{\pi}-\pi^{*}\|$, which is the error between our P-PGDPO policy $\widehat{\pi}$ and the true optimal policy $\pi^{*}$. The bound depends on two primary sources of error. The first is the FOC residual, denoted by $\varepsilon$, which measures how much the policy from the warm-up stage violates the optimality conditions of PMP. The second is the BPTT estimation error, $\delta_{\mathrm{BPTT}}$, in the costate estimate $\widehat{\lambda}$. This estimation error itself stems from two numerical sources: the time discretization of the SDEs (with time step $\Delta t$) and the Monte Carlo sampling used for the expectation (with batch size $M$). The following theorem encapsulates this relationship, with full technical details provided in Appendix C.
{restatable}
[Policy Gap Bound]theoremthm:policy_gap Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$. Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
where:
  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
The proof of Theorem 1, detailed in Appendix C, quantifies the intuition discussed above. The logical chain proceeds as follows: a small FOC-gap ($\varepsilon$) ensures that the policy being evaluated is “close” to satisfying the PDE’s structural requirements. Due to the smoothing property of the uniformly parabolic PDE—guaranteed by regularity theory in Sobolev spaces (e.g., Dong and Kim, (2009))—this “closeness” is not confined to the value function alone but propagates to its derivatives. With the gradients (costates) being verifiably close to the optimal ones, the final step of projecting them through the PMP map becomes a well-posed, stable operation, leading to a policy that is also close to the true optimum.
Our formal analysis demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin FOC gap ($\varepsilon$) from the warm-up stage guarantees that the P-PGDPO policy $\widehat{\pi}$ is provably close to the true optimum $\pi^{*}$, up to numerical errors from simulation. The bound explicitly shows that to achieve a highly accurate policy, one must not only ensure the warm-up is effective (small $\varepsilon$) but also use a sufficiently small time step ($\Delta t$) and a large batch size ($M$) to control the BPTT estimation error. When the objective function is relatively flat near the optimum, the FOC residual $\varepsilon$ for the warm-up policy tends to be small, making the numerical error terms the primary drivers of the final policy gap.
This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these numerically-controlled and rapidly stabilizing costate estimates into the analytical PMP formulas yields a near-optimal policy.
@@
Despite these formal hurdles, the fundamental algorithmic rationale of P-PGDPO remains powerfully persuasive for broader applicability. The core innovation is the decoupling of costate estimation from direct policy construction. Standard end-to-end methods often conflate these two tasks, making it difficult to learn the policy’s fine structure in a ‘flat’ optimization landscape. P-PGDPO, by contrast, leverages BPTT for what it excels at—efficient gradient (costate) estimation—and then applies the robust, analytical structure of PMP to deterministically map these estimates to a high-fidelity policy. This principle, strongly supported by our numerical results, suggests that P-PGDPO is a practical and powerful tool for a wider class of challenging continuous-time control problems, even in cases where the conditions required for the proof may not strictly hold.